• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lhuillier

    Log entry time 16:34:15 on June12,2004

    Entry number 125458

    keyword=Day shift summary

    Started with "production" run at 30uA.
    Then we ramped up the beam current by steps of 2uA. The right septum tripped at 35uA. The temperature of the upper coil was 7.80K just before the trip.
    We continued studies on the left septum. Changing the raster size didn't affect the coil temperature.
    Then Jay took ~2h of beam to finish his test plan for beam tune.
    End of the shift was dedicated to further studies of septum heating. Goal is to see the sensitivity of the septum heating to beam position on target. This test is still going on at the end of the shift.